Commercial Network Services
Network Issues
You are here: Portal Home > Network Issues > Resolved

Test Status Last Upd Next Upd Upd Interval Jan Uptime Dec Uptime
Commercial Network Services (LA) Ping
UP
05:38 AM 05:43 AM 5 min 100.000% 100.000%
Commercial Network Services (NYC) Ping
UP
05:38 AM 05:43 AM 5 min 100.000% 100.000%
Commercial Network Services (UK) Ping
UP
05:38 AM 05:43 AM 5 min 100.000% 100.000%

Website monitoring by Hyperspin

NYC Migration Saturday (Resolved)
Affecting Other - All NY Hosted Services | Priority - High

We will be moving equipment to a new and larger location within the NY datacenter over the next couple Saturdays.  During this time services may be unavailable as equipment is physically moved.  We will resume the services as soon as possible, as they come back online in the new location.
 
No reboots are expected, but services will be unavaiable during the actual move of the respective machine(s).

UPDATE:  Most all services in NY are back online.  We are working on a couple of remaining items to bring everything up as fast as possible.

All work is complete.  During this move, some backend equipment was also upgraded with new bonded 10Gb uplinks.  enjoy






Date - 12/09/2017 01:02 - 12/10/2017 12:04
Last Updated - 12/10/2017 12:04

Network Maintenance NYC (Resolved)
Affecting Other - All NY hosted systems | Priority - High

We are completing maintenance on our network in the NY/NJ region.  There may be brief periods of no connectivity while this work is completed.

Date - 12/02/2017 10:46 - 12/02/2017 13:20
Last Updated - 12/02/2017 10:46

Level3 Transit (Resolved)
Affecting Other - Subscribers using Level 3 transit | Priority - High

We are investigating a possible service issue with Level 3 in North America.  This issue is affecting subscribers with Internet access providers who use L3 with no other alternatives.

The issue is creating lag and disconnections between subscribers and their hosted CNS service.

More information will be posted as soon as it is available.


UPDATE:  Seems to have stabilized.  We will continue to monitor and open this alert again if the issue returns.





Date - 11/27/2017 09:59
Last Updated - 11/27/2017 11:33

Power issue in NY (Resolved)
Affecting Server - VMM | Priority - Low

We are working to resolve a power distribution issue affecting a limited number of VMs in NY.  The issue is causing CPU modulation, which results in a slow server.

UPDATE:  Engineers are setting up temporary power to bring back full performance while damaged equipment is replaced.

INFO:  A power issue destroyed a PDU, network switch and multiple PSU's in about 5 servers.  The network in this rack switched to a redundant switch and the damaged switch has already been returned to normal service.  We have completed running new (temporary) power and replacing PSU's to restore normal quality of service while a proper repair is completed.

Please let CNS Support know if you are experiencing any issues.

As QoS has been restored, priority of this issue has been reduced.  We do not anticipate any further impact to services.

UPDATE:  Full redundancy restored.  All systems normal






Date - 09/04/2017 10:40
Last Updated - 09/06/2017 12:45

Peering to AMS-IX (Resolved)
Affecting Other - Peering to AMS-IX | Priority - Medium

Peering to AMS-IX has flapped.  This will cause brief connectivity issue for traffic that transits the connection while routes reconverge.  We are investigating the cause.  We will update this notice with more information as soon as a review is complete.

---

The original notice listed the IX as DE-CIX.  It is actually AMS-IX

---

We are observing continued instability in the circuit and so the port has been administratively shutdown until the issue is resolved.  Traffic will route around the IX.  No major impact is expected.

---

The problem circuit has been repared and the network has been normalized.

The circuit vendor has responded they are experiencing a major outage.  We will leave the IX offline until they call all clear.



Date - 05/15/2017 07:32
Last Updated - 05/17/2017 07:44

Amsterdam P2P down (Resolved)
Affecting Other - Peering links into Amsterdam | Priority - Medium

Service to Amsterdam has been restored.  The network has been normalized.
____________

A dedicated circuit connecting our network to an IX in Amsterdam has gone down.  Possible fiber cut.  We have raised a ticket with the vendor.


This may cause increased latency to parts of Europe and Africa, but it will not create an outage because there are multiple redundant routes to the affected networks.

We will update this notice as soon as service to Amsterdam IX is restored.




Date - 05/04/2017 08:55
Last Updated - 05/04/2017 21:35

FXCM routing thru China Mobile (Resolved)
Affecting Other - FXCM routing | Priority - Critical

We have received a report of high latency to FXCM.  Our investigation has revealed that China Mobile is advertising their prefix, causing this traffic to route through their network.

We are working to filter the advertisement.

UPDATE:  We have filtered out the hyjacked route.  All appears normal.

We will continue to monitor.



Date - 01/29/2017 18:53 - 01/29/2017 23:15
Last Updated - 01/29/2017 19:07

Subscriber Control Panel Upgrade (Resolved)
Affecting Other - Subscriber Control Panel | Priority - Low

Please be advised the CNS Subscriber Control Panel will be unavailable during scheduled maintenance so we can install software upgrades.  If you require assistance during this time, please call the helpdesk at any number below:

San Diego, CA: +1 (619) 225-7882

Los Angeles, CA: +1 (213) 769-1787

New York, NY: +1 (646) 930-7435

London, UK: +44 (2035) 191453

 

 

Date - 07/02/2016 00:30 - 07/04/2016 07:43
Last Updated - 06/30/2016 15:51

Circuit issue in NY region (Resolved)
Affecting Other - NY POP | Priority - High

Two circuits in the NY region have been cut somewhere outside our cage.  We are working with the datacenter to get them re-run as fast as possible.  In the mean time, increased latency may be experienced as we have routed around the trouble.

One of two have been replaced.  Currently waiting on second line run and we will then normalize the network.

Damaged fiber has been replaced.  We are now testing before normalizing the network.

All tests have passed with no packet loss.  We are proceeding to normalize the network.

Network has been normalized.  We are observing to confirm.

Some routes are still reconverging.  we are observing












Date - 06/13/2016 08:03 - 06/14/2016 07:57
Last Updated - 06/13/2016 12:05

Network Maintenance [extended] (Resolved)
Affecting Other - Network | Priority - Low

We are performing maintenance on our network.  During this time some routes may be intermittent.  We are working to complete the maintenance as fast as possible.

Hardware issue during maintenance created complications.  We are finishing it up.  Apologizes for the delay.


Date - 06/12/2016 00:18 - 06/12/2016 18:20
Last Updated - 06/12/2016 10:37

« Prev Page   Next Page »